# Class-Discriminative Attention Maps for Vision Transformers

### v2

@@ 4.2.3 Evaluation of compactness
Sparsity and shrinkage are evaluated by our compactness evaluation, which is one of the major desired properties in explainability of deep learning [44]. While sparsity may not always imply the most accurate estimation at a single data point, a bias-variance trade-off is well-known in machine learning [47]. Ideally, compact explanations (e.g., sparse and shrunken importance estimators) are preferred, when other properties such as correctness and class sensitivity are constant.
@@ 5 Application to Medical Images
To demonstrate CDAM for another use case we apply it to lung nodule malignancy and biomarker prediction using the LIDC dataset [23].
@@ 5.1 Malignancy Prediction
After preprocessing, the LIDC data consists of 443 benign and 411 malignant lung CT scans. Training, validation, and test sets (in the ratios of 0.7225, 0.1275, 0.15) were stratified by and balanced according to these labels, e.g., benign and malignant. We fine-tuned a ViT model, with a DINO backbone pretrained on the ImageNet, for 50 epochs. In a parameter sweep, we varied the number of trainable layers ($10-50$) and dropout rates ($0.0-0.09$), where the learning rate was exponentially decaying ($\alpha=0.0003$ and $\beta=0.95$). The best accuracy on the test set of 0.85 was obtained with 50 trainable layers and dropout rate of 0.003222It might be possible to increase the model performance via further investigation. Here, we use a straightforward model and training scheme to demonstrate the application of CDAM on medical images.. Attention maps and CDAM are obtained based on this model.
The obtained attention maps suggest that the model focuses on patches with nodule fragments (Fig. 12(b)). However, we note that attention maps almost always focus on pulmonary nodules without taking into account the downstream classification into benign and malignant samples. CDAM provides detailed structures where positive and negative values are indicated by orange and blue, respectively. In this binary classification where 0 is for benign and 1 for malignancy, input pixels with positive values in CDAMs are driving the classification towards malignancy.
@@ 5.2 Biomarker prediction
We turn our attention to clinical biomarkers in LIDC that are routinely used by medical practitioners. Inspired by the concept bottleneck model (CBM) [48], we fine-tuned a ViT model pre-trained with DINO on 8 biomarkers (subtlety, calcification, sphericity, margin, lobulation, spiculation, texture, and diameter). Essentially, a regression model was built on a pre-trained ViT model with the mean squared error (MSE) as loss function. For investigation of LIDC and incorporation of CBM, see [49].
@@ 6 Discussion
Class-Discriminative Attention Map (CDAM) is a gradient-based extension of attention maps, that provides strong discrimination with respect to a chosen class or concept, while exhibiting excellent fidelity and compactness. We introduce additional variants of CDAM, inspired by SmoothGrad ([15]) and IntGrad [16]), which improve the performance on the quantitative benchmarks. Our proposed methods retain appealing characteristics of attention maps, namely high-quality semantic segmentation in self-supervised ViTs. The proposed importance estimation for CDAM scales the attention scores by how relevant they are for a target class or concept (Section A.1). Therefore, zero attention scores remain zero in CDAM and it produces compact explanations with high sparsity and shrinkage.
@@
Post-hoc explanation methods for DNNs aim to make the full decision-making process of the model more transparent by providing an approximation with certain desirable properties, such as correctness, sensitivity, and compactness. Particularly, many explanation methods that estimate feature importance are not very useful because they give almost identical results when targeting different classes in the model output [27, 38]. Even randomization in the model weights is often shown to have minimal impact on explanations (i.e., saliency maps [38]). We find, both qualitatively and quantitatively, that CDAM is highly sensitive to the targeted class, assigning positive importance scores to objects corresponding to the target class and negative ones to other (semantically distinct) objects in the image.
Besides explaining predictions of the classifier on the top of ViT, we introduce several variants. Firstly, by probing the latent space of the ViT (Fig. 4), self-supervised CDAM provides importance scores specific to the user-defined concept. In the context of SSL, where class labels are absent, this makes it a valuable technique to investigate the representations learned by the model, as shared by user-selected samples. To some degree, this can alternatively be done by looking at the attention maps of the different heads, which tend to specialize in different categories of objects or parts thereof [9]. This approach is, however, limited to the categories that the attention heads happened to have learned. Secondly, we introduce Smooth and Integrated CDAM, which essentially average a series of (vanilla) CDAMs. Notably, instead of adding noise to or modifying the input images as in SmoothGrad [15]) or IntGrad [16], respectively, our methods act on tokens in the final transformer block.
Our application on medical images examplifies the need for fine-grained class-sensitive explanations offered by CDAM. Other explainability methods including attention maps are often highlighting whole nodules and tumors. But they are uninformative about the innerworkings of the ViT model predicting malignancy or biomarkers. While medical implication of these explanations are beyond the scope of this paper, the proof of concept application demonstrates usability of CDAM in cancer images. As the methods of explainable AI (XAI) becomes incorporated into practices, we foresee a greater level of human-computer interaction where class sensitivity will play a critical role for utility of explanations.
Understanding the decision-making process of deep neural networks is intrinsically challenging due to the vast parameter space and non-linearity, which lead to complex interactions among the patch tokens in ViT. Post-hoc explanation methods for DNNs approximate the full operating characteristics of the model with certain desirable properties. Our results suggest that CDAM is capable of achieving a high degree of correctness, contrastivity, and compactness. Note that CDAM is not explicitly designed to satisfy these criteria. For example, we could impose further sparsity constraints that would increase compactness. However, over-optimizing an explanation method for certain properties could have undesirable effects [^3] Goodhart’s Law states that “when a measure becomes a target, it ceases to be a good measure” [50].. While this paper focuses on a ViT model, CDAM is applicable more broadly to all transformer models with self-attention, including large language models (LLMs) [51].
Acknowledgments
This work was funded by the ERA-Net CHIST-ERA grant [CHIST-ERA-19-XAI-007] long-term challenges in ICT project INFORM (ID: 93603), by the National Science Centre (NCN) of Poland [2020/02/Y/ST6/00071]. This research was carried out with the support of the Interdisciplinary Centre for Mathematical and Computational Modelling University of Warsaw (ICM UW) under computational allocation no GDM-3540; the NVIDIA Corporation’s GPU grant; and the Google Cloud Research Innovators program.
@@ A.1 Proportionality of CDAM to attention maps
The proposed class-discriminative attention scores scale the attention scores. Empirically, in all of our evaluation tasks and applications (e.g., Fig. 11), we observe that features with attention scores $A_{i}=\text{softmax}\left(\frac{Q_{\texttt{[CLS]}}K^{\top}_{i}}{\sqrt{d_{k}}}\right)$ of 0 have class-discriminative attention scores of 0:
@@
$S_{i,c}$ $=\sum_{j}T_{ij}\frac{\partial g(l,l_{c})}{\partial T_{ij}}$
$=\sum_{j}T_{ij}\frac{\partial}{\partial T_{ij}}\sum_{k}l_{k}l_{c,k}$
$=\sum_{j}T_{ij}\frac{\partial}{\partial T_{ij}}\sum_{k}h_{k}\left(\sum_{m}A_{m}V_{m}+\texttt{[CLS]}\right)l_{c,k}$
$=\sum_{j}T_{ij}\sum_{k,m,n}\nabla_{n}h_{k}\times\left(\frac{\partial A_{m}}{\partial T_{ij}}V_{mn}+A_{m}\frac{\partial V_{mn}}{\partial T_{ij}}\right)l_{c,k}$
$=\sum_{j,k,m,n}T_{ij}\nabla_{n}h_{k}\frac{\partial A_{m}}{\partial T_{ij}}V_{mn}l_{c,k}+\sum_{k,n}A_{i}V_{in}\nabla_{n}h_{k}l_{c,k},$ (10)
@@
Eq. 9 implies that in Section A.1, all terms that are not proportional to $A_{i}$ are zero or cancelled out and we are left with
@@
In the special case that $h$ is the identity function, we obtain an even simpler interpretation of CDAM due to $\nabla_{n}h_{k}(f)=\partial_{f_{n}}f_{k}=\delta_{nk}$. Assuming $\frac{\partial A_{m}}{\partial T_{ij}}=0$, Section A.1 simplifies to $S_{i,c}=A_{i}V_{i}\cdot l_{c}$, which means that $A_{i}$ is scaled by the similarity of $V_{i}$ and $l_{c}$.
We note that introducing $h$ attention heads does not change the above derivations qualitatively. Instead of the sum $Z=\sum_{m}A_{m}V_{m}$, there will be a term $\text{Concat}(Z^{1},Z^{2},..,Z^{h})W^{O},W^{O}\in\mathbb{R}^{hd_{v}\times d_{\text{model}}}$ so that in the case of multi-head attention the equivalent of Section A.1 reads
@@
and since $A^{n}_{i}>0,A_{i}=\sum_{n}\frac{A_{i}^{n}}{h}$ implies that $A^{i}=0\Rightarrow A_{i}^{n}=0,\forall n$ we can again conclude that in order for Eq. 9 to hold all terms that are not proportional to $A_{i}$ either vanish or cancel out. For multi-head attention we thus arrive at
@@
Although our observation $A_{i}=0\Rightarrow S_{i,c}=0$ implies that all terms that are not proportional to $A_{i}$ in Section A.1 should vanish, we don’t have a theoretical argument why that would be the case. It seems unlikely that the various terms in the first sum cancel out, which leaves $\frac{\partial A_{m}}{\partial T_{ij}}=0$ as an explanation. However, in general, $A_{i}$ depends on $T_{i}$ through $K_{i}$ and one would therefore expect the derivative to be non-zero.
@@
$S_{i,c}$ $=\sum_{j}T_{ij}\frac{\partial f_{c}}{\partial T_{ij}}$
$=\sum_{j}T_{ij}\frac{\partial}{\partial T_{ij}}\sum_{k}l_{k}W^{C}_{kc}$
$=\sum_{j,k,m,n}T_{ij}\nabla_{n}h_{k}\frac{\partial A_{m}}{\partial T_{ij}}V_{mn}W^{C}_{kc}+\sum_{k,n}A_{i}V_{in}\nabla_{n}h_{k}W^{C}_{kc},$ (14)
@@
A.2 Results for a random subset of ImageNet
Figure S1: Box sensitivity curves obtained from the random ImageNet subset
Figure S2: MIF/LIF perturbation curves obtained from the random ImageNet subset.
@@ A.2 Results for a random subset of ImageNet
A.3 Examples of CDAM for a known class on the ImageNet
Figure S3: Examples for CDAM and other explanation methods on the ImageNet. The indicated class is the one predicted by the model and is the target for the explanation methods.
Figure S4: Examples for CDAM and other explanation methods on the ImageNet. The indicated class is the one predicted by the model and is the target for the explanation methods.
Figure S5: Examples for CDAM and other explanation methods on the ImageNet. The indicated class is the one predicted by the model and is the target for the explanation methods.
A.4 Examples for CDAM on the Lung Image Database Consortium (LIDC)
Figure S6: Examples for the Attention Map and CDAM based on a ViT malignancy model trained on the lung nodule data from LIDC.

### v3

@@ 4.2.3 Evaluation of compactness
[…]
CDAM 0.88
Integrated CDAM 0.88
Smooth CDAM 0.75
Relevance Propagation 0.48
Partial LRP 0.74
Attention map 0.56
SmoothGrad 0.8
IntGrad 0.84
GradCAM 0.54
Input $\times$ Grad 0.88
Table 3: Compactness, measuring sparsity and shrinkage, is defined by the fraction of pixels with an importance score lower or equal to 5% of the maximum score, evaluated on the $X_{\text{multi}}$ dataset.
Sparsity and shrinkage are evaluated by our compactness evaluation, which is one of the major desired properties in the explainability of deep learning (Nauta et al., 2023a). While sparsity may not always imply the most accurate estimation at a single data point, a bias-variance trade-off is well-known in machine learning (Lahlou Kitane, 2022). Ideally, compact explanations (e.g., sparse and shrunken importance estimators) are preferred, when other properties such as correctness and class sensitivity are constant.
@@ 5 Application to Medical Images
To demonstrate CDAM for another use case we apply it to nodule malignancy and biomarker prediction using the Lung CT scans in the LIDC dataset (Armato III et al., 2011). Here, we show CDAMs from DINO-based models, we also showcase use of DeIT (Dosovitskiy et al., 2020; Touvron et al., 2021) and DINOv2 (Oquab et al., 2024; Darcet et al., 2024) backbones that are fine-tuned on the LIDC dataset in Section A.5.2.
@@ 5.1 Malignancy Prediction
After preprocessing, the LIDC data consists of 443 benign and 411 malignant lung CT scans. Training, validation, and test sets (in the ratios of 0.7225, 0.1275, 0.15) were stratified by and balanced according to these labels, e.g., benign and malignant. We fine-tuned a ViT model, with a DINO backbone pre-trained on the ImageNet, for 50 epochs. In a parameter sweep, we varied the number of trainable layers ($10-50$) and dropout rates ($0.0-0.09$), where the learning rate was exponentially decaying ($\alpha=0.0003$ and $\beta=0.95$). The best accuracy on the test set of 0.85 was obtained with 50 trainable layers and dropout rate of 0.003111It might be possible to increase the model performance via further investigation. Here, we use a straightforward model and training scheme to demonstrate the application of CDAM on medical images.. Attention maps and CDAM are obtained based on this model.
The obtained attention maps suggest that the model focuses on patches with nodule fragments (Fig. S11(b)). However, we note that attention maps almost always focus on pulmonary nodules without taking into account the downstream classification into benign and malignant samples. CDAM provides detailed structures where positive and negative values are indicated by orange and blue, respectively. In this binary classification where 0 is for benign and 1 for malignancy, input pixels with positive values in CDAMs are driving the classification towards malignancy.
@@ 5.2 Biomarker prediction
We turn our attention to clinical biomarkers in LIDC that are routinely used by medical practitioners. Inspired by the concept bottleneck model (CBM) (Koh et al., 2020), we fine-tuned a ViT model pre-trained with DINO on 8 biomarkers (subtlety, calcification, sphericity, margin, lobulation, spiculation, texture, and diameter). Essentially, a regression model was built on a pre-trained ViT model with the mean squared error (MSE) as loss function. For investigation of LIDC and incorporation of CBM, see (Brocki & Chung, 2023a).
@@
Figure 7: Explanations for the biomarker prediction model. Eight biomarkers available in the LIDC dataset are used to fine-tune a ViT regression model with a DINO backbone. CDAMs provide distinct and relevant explanations for different biomarkers. For more examples, see Section A.4.
@@ 6 Discussion
Class-Discriminative Attention Map (CDAM) is a gradient-based extension of attention maps, that provides strong discrimination with respect to a chosen class or concept, while exhibiting excellent fidelity and compactness. Our proposed methods retain appealing characteristics of attention maps, namely their high-quality semantic segmentation. The proposed importance estimation for CDAM scales the attention scores by how relevant they are for a target class or concept (Section A.1). Therefore, zero attention scores remain zero in CDAM and it produces compact explanations with high sparsity and shrinkage. Besides explaining predictions of the classifier on the top of ViT, CDAM can also provide importance scores specific to the user-defined concept. In the context of self-supervised models, where class labels are absent, this makes it a valuable technique to investigate the latent representations learned by the model.
@@
Post-hoc explanation methods for DNNs aim to make the full decision-making process of the model more transparent by providing an approximation with certain desirable properties, such as correctness, sensitivity, and compactness. Particularly, many explanation methods that estimate feature importance are not very useful because they give almost identical results when targeting different classes in the model output (Rudin, 2019; Adebayo et al., 2018). Even randomization in the model weights is often shown to have minimal impact on explanations (i.e., saliency maps (Adebayo et al., 2018)). We find, both qualitatively and quantitatively, that CDAM is highly sensitive to the targeted class, assigning positive importance scores to objects corresponding to the target class and negative ones to other (semantically distinct) objects in the image.
We introduce Smooth and Integrated CDAM, which essentially average a series of (vanilla) CDAMs. Notably, instead of adding noise to or modifying the input images as in SmoothGrad (Smilkov et al., 2017) or IntGrad (Sundararajan et al., 2017), respectively, our methods act on tokens in the final transformer block. While we conducted a number of quantitative and qualitative evaluations, there is no single XAI method that triumph in all scenarios. However, considering that Smooth and Integrated CDAM require computing $n$ vanilla CDAMs (e.g., $n=50$), we recommend first trying vanilla CDAMs. If high fidelity is singularly desirable, we suggest utilizing Smooth CDAM.
Our application on medical images exemplifies the need for fine-grained class-sensitive explanations offered by CDAM. Other explainability methods including attention maps often highlight whole nodules and tumors. However, they are uninformative about the inner workings of the ViT model predicting malignancy or biomarkers. As the methods of explainable AI (XAI) are considered for practical implementations, the investigation of human interactions with different explainability methods represents a promising direction for future studies.
Broader Impact Statement
The introduction of our importance estimators would enhance the transparency, trustworthiness and accountability of vision transformer models. By providing clearer insights into model decision-making processes, accurate explainability methods help foster broader adoption of AI technologies. However, the explanations could be manipulated by adversarial input perturbation or data poisoning. In turn, enhanced explanations may inadvertently expose vulnerabilities, making models more susceptible to adversarial attacks or malicious exploitation. Therefore, it is imperative to continuously monitor and ensure the robustness of explanation methods while implementing safeguards to mitigate the risks associated with adversarial attacks.
@@ A.1 Proportionality of CDAM to attention maps
The proposed class-discriminative attention scores scale the attention scores. Empirically, in all of our evaluation tasks and applications (e.g., Fig. S1), we observe that features with attention scores $A_{i}=\text{softmax}\left(\frac{Q_{\texttt{[CLS]}}K^{\top}_{i}}{\sqrt{d_{k}}}\right)$ of 0 have class-discriminative attention scores of 0:
@@
$S_{i,c}$ $=\sum_{j}T_{ij}\frac{\partial g(l,l_{c})}{\partial T_{ij}}$ (10)
$=\sum_{j}T_{ij}\frac{\partial}{\partial T_{ij}}\sum_{k}l_{k}l_{c,k}$ (11)
$=\sum_{j}T_{ij}\frac{\partial}{\partial T_{ij}}\sum_{k}h_{k}\left(\sum_{m}A_{m}V_{m}+\texttt{[CLS]}\right)l_{c,k}$ (12)
$=\sum_{j}T_{ij}\sum_{k,m,n}\nabla^{n}h_{k}\times\left(\frac{\partial A_{m}}{\partial T_{ij}}V_{mn}+A_{m}\frac{\partial V_{mn}}{\partial T_{ij}}\right)l_{c,k}$ (13)
$=\sum_{j,k,m,n}T_{ij}\nabla^{n}h_{k}\frac{\partial A_{m}}{\partial T_{ij}}V_{mn}l_{c,k}+\sum_{k,n}A_{i}V_{in}\nabla^{n}h_{k}l_{c,k},$ (14)
@@
Eq. 9 implies that in Eq. 10, all terms that are not proportional to $A_{i}$ are zero or cancelled out and we are left with
@@
In the special case that $h$ is the identity function, we obtain an even simpler interpretation of CDAM due to $\nabla_{n}h_{k}(f)=\partial_{f_{n}}f_{k}=\delta_{nk}$. Assuming $\frac{\partial A_{m}}{\partial T_{ij}}=0$, Eq. 10 simplifies to $S_{i,c}=A_{i}V_{i}\cdot l_{c}$, which means that $A_{i}$ is scaled by the similarity of $V_{i}$ and $l_{c}$.
We note that introducing $h$ attention heads does not change the above derivations qualitatively. Instead of the sum $Z=\sum_{m}A_{m}V_{m}$, there will be a term $\text{Concat}(Z^{1},Z^{2},..,Z^{h})W^{O},W^{O}\in\mathbb{R}^{hd_{v}\times d_{\text{model}}}$ so that in the case of multi-head attention the equivalent of Eq. 10 reads
@@
and since $A^{n}_{i}>0,A_{i}=\sum_{n}\frac{A_{i}^{n}}{h}$ implies that $A^{i}=0\Rightarrow A_{i}^{n}=0,\forall n$ we can again conclude that in order for Eq. 9 to hold all terms that are not proportional to $A_{i}$ either vanish or cancel out. For multi-head attention, we thus arrive at
@@
Although our observation $A_{i}=0\Rightarrow S_{i,c}=0$ implies that all terms that are not proportional to $A_{i}$ in Eq. 10 should vanish, we don’t have a theoretical argument why that would be the case. It seems unlikely that the various terms in the first sum cancel out, which leaves $\frac{\partial A_{m}}{\partial T_{ij}}=0$ as an explanation. However, in general, $A_{i}$ depends on $T_{i}$ through $K_{i}$ and one would therefore expect the derivative to be non-zero.
@@
$S_{i,c}$ $=\sum_{j}T_{ij}\frac{\partial f_{c}}{\partial T_{ij}}$ (18)
$=\sum_{j}T_{ij}\frac{\partial}{\partial T_{ij}}\sum_{k}l_{k}W^{C}_{kc}$ (19)
$=\sum_{j,k,m,n}T_{ij}\nabla^{n}h_{k}\frac{\partial A_{m}}{\partial T_{ij}}V_{mn}W^{C}_{kc}+\sum_{k,n}A_{i}V_{in}\nabla^{n}h_{k}W^{C}_{kc},$ (20)
@@
A.2 Quantitative Evaluations
Figure S1: Scatter plot of raw attention scores and token importance scores obtained from 200 randomly selected samples of the ImageNet validation set $X_{\text{val}}$.
Figure S2: Fidelity of importance estimators based on MIF/LIF perturbation curves obtained from the $X_{\text{multi}}$ dataset. The change in the model output is measured as a proportion of input features are perturbed according to the Least Important First (LIF) or the Most Important First (MIF) orders. Importance estimators are evaluated by the area between LIF and MIF perturbation curves, which is equivalent to the area underneath $f_{\text{LIF}}$-$f_{\text{MIF}}$ (right). See Table 3.
Figure S3: Correctness of importance estimators as measured by box sensitivity evaluation. The change in the model output (logit) is measured as pixels within a $s\times s$ box are perturbed. The sum of their importance scores would correlate with the change in the model output, if importance estimators are accurate. This is repeated 100 times per sample and the Pearson correlation between the sum of importance scores and the change in the model output is calculated, based on the $X_{\text{multi}}$ dataset.
Figure S4: MIF/LIF perturbation curves for the evaluation of class-discrimination. Different classes are targeted for obtaining importance scores and measuring model output (Section 3.5), a large $\Delta A_{\text{box}}$, therefore, indicates importance scores that are class-discriminative. See Table 2
Figure S5: Box sensitivity curves for the evaluation of class-sensitivity
Figure S6: Box sensitivity curves obtained from the random ImageNet subset $X_{\text{val}}$. Refer to Section 3.5 for data description.
Figure S7: MIF/LIF perturbation curves obtained from the random ImageNet subset $X_{\text{val}}$. Refer to Section 3.5 for data description.
@@ A.2 Quantitative Evaluations
A.3 Examples of CDAMs from the ImageNet
Figure S8: Additional examples for CDAM and other explanation methods on the ImageNet. The indicated class is the one predicted by the model and is the target for the explanation methods.
Figure S9: Additional examples for CDAM and other explanation methods on the ImageNet. The indicated class is the one predicted by the model and is the target for the explanation methods.
Figure S10: Additional examples for CDAM and other explanation methods on the ImageNet. The indicated class is the one predicted by the model and is the target for the explanation methods.
A.4 Examples for CDAMs from lung CT scans
Figure S11: Explanations for the malignancy prediction model. A ViT-based classifier fine-tuned on the lung CT scans (LIDC) was used to predict benign vs. malignant lung nodules, followed by obtaining their attention maps and CDAMs. Orange and blue correspond to positive and negative values, respectively.
@@ A.4 Examples for CDAMs from lung CT scans
A.5 Additional ViT architectures and training strategies
While the main manuscript has used ViT models pre-trained using DINO (Caron et al., 2021), other architectures and pre-trained weights can be used for CDAM. In this section, we show examples of alternative ViT models that are trained in alternative manners. As expected, pre-trained ViT models that produce high-quality attention maps result in high-quality CDAMs. In particular, note that the patch size of DINO (Caron et al., 2021) is $8\times 8$ compared to much larger patch sizes used in other backbones.
Table S2: Pre-trained ViT models and training strategies.
Name Model sizes Patch sizes References
DeIT (original ViT) 86M 16 Dosovitskiy et al. (2020); Touvron et al. (2021)
SWAG 86M 16 Singh et al. (2022)
DINOv2 (with registers) 21M, 86M 14 Oquab et al. (2024); Darcet et al. (2024)
DINO (used in the main text) 21M, 85M 8, 16 Caron et al. (2021)
A.5.1 Alternative ViT models for the ImageNet
We use a ViT model pre-trained on the ImageNet-1K dataset using supervised weakly through hashtags (SWAG) (Singh et al., 2022). Unlike DINO and other self-supervised ViT models, the SWAG model has been trained end-to-end. Therefore, without pre-training, we can obtain attention maps and CDAMs directly. Three examples are shown in Fig. S14.
Figure S14: Attention maps and CDAMs from the ImageNet-1K dataset, based on the pre-trained ViT model backbone using SWAG Singh et al. (2022).
In general, ViT models require a massive amount of training data and therefore often trained in self-supervised manner. As shown in DINO (Caron et al., 2021), self-supervised learning can not only improve performance of fine-tuned models, but also improve attention maps. CDAM relies on attention maps, therefore the quality of attention maps strongly influence that of CDAM. Note that the SWAG model has a patch size of $16\times 16$, which results in lower-resolution attention maps and CDAMs.
A.5.2 Alternative ViT models for the LIDC
We demonstrate the use of CDAM with ViT models trained using Data-efficient image Transformers (DeIT) (Touvron et al., 2021) and DINOv2 (Oquab et al., 2024). Note that despite the use of the same acronym, DINOv2 is substantially different from DINO (Caron et al., 2021) due to using a different training strategy introduced in iBOT (Zhou et al., 2022) on a much larger scale (Oquab et al., 2024). The original DINO (Caron et al., 2021) was pre-trained on the ImageNet, whereas DINOv2 (Oquab et al., 2024) used a newly curated dataset LVD-142M. Additional tokens called “registers” were introduced to improve attention maps of DINOv2 (Darcet et al., 2024).
The LIDC dataset was split into 5 folds and stratified according to malignancy status. The best-performing models were chosen based on means of Accuracy (ACC) and Mean Squared Error (MSE) over 5 folds, in the validation set. In a parameter sweep, we varied the number of trainable layers (10 - all), dropout rates in the backbone (0.0 - 0.12), batch size (8 - 32) and learning rate scheduler parameters. The learning rate was scheduled with CyclicLR scheduler. Binary Cross Entropy Loss and Huber Loss were used to train classification and regression models, respectively. Weights optimization was performed with Adam optimizer and random rotation was applied as data augmentation. For malignancy classification, the best-performing models based on DeIT and DINOv2 achieved a mean ACC of 0.896 and 0.904, respectively. For biomarker regression, the best-performing models based on DeIT and DINOv2 archived the mean MSE of 0.409 and 0.35, respectively.
Figure S15: Attention Maps and CDAMs for malignancy models using ViT backbones pre-trained with DeIT and DINOv2.
Figure S16: Attention Maps and CDAMs for biomarkers models using ViT backbones pre-trained with DeIT
Figure S17: Attention Maps and CDAMs for biomarkers models using ViT backbones pre-trained with DINOv2